Agent Fox Mulder wants to form a XY Files portfolio consisting of 70% Stock X and 30% Stock Y. What is the XY Files portfolio expected return?

Probability /X/ Y

0.2/ -20%/ 10%

0.6/ 20/ 15

0.2/ 40/ 20

A. 15.7% B. 15.0% C. 21.3% D. 24.0% E. 16.0%
Answer #1

Expected return=Respective return*Respective probability

Expected return of X=(0.2*-20)+(0.6*20)+(0.2*40)=16%

Expected return of Y=(0.2*10)+(0.6*15)+(0.2*20)=15%

Expected return of portfolio=(16*0.7)+(15*0.3)

which is equal to

=15.7%